# Hawking radiation in a charged, cosmological black hole

### v1

@@ Abstract
We study the effective temperature, as a rate of gravitational redshift, of the Hawking modes perceived by a radially free falling observer at an arbitrary location in a Reissner-Nordström-(Anti-)de Sitter (RN(A)dS) spacetime. Observers of both positive and negative specific energies have been taken into account. Particular emphasis is on the behaviour of the modes at the inner horizon, and therein on the validity of the strong cosmic censorship conjecture across the physically permissible parameter space of the RNdS metric. Lastly, the behavior of the adiabatic control function has been analysed over the position space of the observer to determine the regimes where the effective temperature function yields a Planckian spectrum of thermal radiation.
@@ I Introduction
The formulation relevant to this paper is based upon the result that an exponential relation between the affine parameters (u,U) of past and future null infinities results in Hawking radiation [4, 5, 6]. A function $\kappa(\textit{u})$ is defined, dependent on the parametrized relationship U(u), that primarily governs the peeling properties of the null geodesics [7]. It has been shown that given an exponential relation between u and U, and given that the adiabatic condition [Eq. (2)] holds, a thermal flux will be generated, whose temperature is proportional to $\kappa(\textit{u})$. In the limit of a static black hole, $\kappa$ becomes equal to the surface gravity $\varkappa_{+}$ of the event horizon. It is possible to recast the definition of the effective temperature $\kappa$ as the rate of exponential redshift of the Hawking modes. In the case of a collapsing mass, the modes experience a blueshift as they approach the mass and are subject to a gravitational redshift as they climb out. Another explanation that well complements this rather non-intuitive formalism notes that the vacuum state in which the thermal radiance is observed is related to the local Minkowski frame of the observer via an exponential scale transformation [8]. This inherently draws upon the analogous relation between the Unruh effect and the Hawking radiation—an accelerated observer in the Minkowski spacetime, and observer at a constant distance from the black hole will both detect a flux of particles [9].
@@
Of significant consequence in the study of an RNdS black hole, is the existence of a cosmological horizon (at radius $r_{c}$), a boundary that causally separates an observer inside ($r<r_{c}$) from the external universe ($r>r_{c}$). It is characteristically very similar to the event horizon, following the same classical and thermodynamic laws, therein radiating particles with a thermal spectrum [1]. A rigorous study on the thermodynamic properties of the cosmological horizon was presented by Hawking and Gibbons [18], extending to particle creation in de Sitter spaces. While the matter of particle creation in de Sitter space had for long been a topic of study, it had always been viewed in an observer-independent frame. If particle creation were observer independent, and the invariance of the de Sitter group [19, 20, 21] were to hold, an observer Lorentz-boosted into a new frame will see exactly the same spectrum as they would have before the boost. This cannot be true unless the rate of particle production was either zero or infinity, zero being the favoured choice of the two. But no particle production would imply no thermal radiation emanating from the cosmological horizon. However it was shown that an observer accelerating in Minkowski spacetime will detect particle production, indicating that particle production must be dependent on the worldline of the observer. An observer moving on a timelike geodesic will detect an isotropic thermal radiation of temperature $T_{H}=\frac{\varkappa_{c}\hbar}{2\pi ck_{B}}$, where $\varkappa_{c}$ is the surface gravity of the cosmological horizon. Once again, the invariance of the de Sitter group requires that any other observer also on a timelike geodesic detect the same radiation—indicating that particle production is in fact observer dependent. We will arrive at the same conclusion, albeit via a different approach. Both the outgoing and the ingoing modes tend to the same isotropic temperature asymptotically far from the black hole, independent of the observer’s specific energy. We will further extend the analysis to negative observer energies and a negative cosmological constant.
@@ II Formalism
II.1 The adiabatic approximation
Consider an emitter with a position characterized by null coordinate $U$, and an observer following a timelike geodesic characterized by null coordinate $u$. The worldlines of the emitter and the observer are connected by a null ray following a geodesic path given by $U(u)$. Without loss of generality, $U(u)$ can be described via an exponential relationship [4], dependent on a function $\kappa(u)$:
The adiabatic condition requires that around any given point $u_{*}$, the value of the function $\kappa_{*}\equiv\kappa(u_{*})$ remains approximately constant. Mathematically, the constraint can be given as
$\left|\frac{1}{\kappa_{*}^{2}}\frac{d\kappa_{*}}{du}\right|\ll 1.$ (2)
$T_{H}=\frac{\kappa(u_{*})}{2\pi}.$ (3)
Having scaled the constants, the effective temperature at any point can be well given by $\kappa(u_{*})$. Of particular note is that this formalism is independent of the position $u_{*}$ of the observer. $\kappa(u)$ by definition, is the parametrized relation between the affine parameters of the emitter and the observer, describing the peeling properties of the null geodesics. It is on this quantity that the Hawking temperature is dependent, rather than the surface gravity itself. In the limit of a static black hole, i.e., as the observer approaches future infinity, $\kappa(u)$ equals the surface gravity, reducing the expression to the well established Hawking temperature. This formulation (3) thereby permits a calculation of the effective temperature perceived by an observer at any point in spacetime.
The exponential approximation was originally derived for an asymptotically flat spacetime. Because the RNdS metric is not asymptotically flat, one might question the validity with which the formalism may be extended to the aforementioned metric. However, it must be reinforced that the given metric represents a static black hole, and does not incorporate a dynamic collapse. The construction of the Unruh vacuum state (Section II.3) permits a modelling of dynamic gravitational collapse in which spacetime becomes asymptotically de Sitter at past null infinity. The derivation of the adiabatic approximation can be extended to spacetimes that are not flat, but de Sitter in the asymptotic past.
Because the observer and the emitter can use their respective proper times to label the affine null parameters, the effective temperature function (1) can be written as
It is this form of the effective temperature function that will be relevant to the subsequent calculations.
II.2 Geodesics in the RNdS Geometry
@@
The temporal component of the four velocity of a radially free-falling observer or emitter is
@@
The components for the four velocity of the observer ($u_{\mu}$) and of the null particle ($k^{\mu}$) are found using the standard conditions
@@
The frequency of the null particle as seen by the observer or emitter is obtained by transforming the wave four vector of the null particle into the frame of the observer or the emitter:
@@
where the upper(lower) signs correspond to outgoing(ingoing) modes respectively.
@@
$E$ is the specific energy of the observer/emitter, which remains constant along the natural geodesic. It was defined earlier as the negative temporal component of the velocity four vector. The value of the specific energy can be easily calculated as the square root of the horizon function at a point where $u_{r}=0$. We can conveniently choose (as we subsequently will) the rest point of the observer to be at the point where the horizon function is at a maximum.
@@
We will define the specific energy to be positive for an observer who starts out from rest at a point between the event and cosmological horizons. An infalling(outfalling) observer may be defined as one who travels to a smaller(larger) spatial radius by virtue of the natural geodesic trajectory. An ingoer(outgoer) is defined by their intended direction of motion. The distinction may not appear to be of great significance in a non-trapping region, where an observer can accelerate enough to be able to oppose the natural geodesic. In a trapping region, such as inside the event horizon, it is not possible to be able to travel to a larger spatial radius, and an attempted outward acceleration will induce the observer into a state of negative energy. An observer beyond the event horizon ($r<r_{+}$), who is infalling and outgoing, will have a specific energy $E<0$. So also, an observer beyond the cosmological horizon who is outfalling but ingoing will also have a negative specific energy. Note that the modes are defined as being either outgoing or ingoing, a state determined by the initial conditions at past null infinity. Outgoing modes inside the event horizon will remain outward directed, but will travel to a smaller spatial radius, eventually undergoing a divergence at the inner horizon.
II.3 Construction of the Unruh State
The RNdS metric, in the form of Eq. (5) depicts an eternal and static black hole. Realistic black holes, however, are not eternal, and are formed via a gravitational collapse. The Unruh quantum state [9] formulates an equivalent to the dynamic collapse by defining a quantum field with the appropriate boundary conditions on the past horizon.
A massless scalar wave equation, in a given metric, can be solved for a set of normal mode solutions $\phi_{\omega lm}$. The radial component of these solutions yields two sets of waves—one coming in from infinity, the other emanating from the past horizon of the black hole. Positive frequencies with respect to a vector $\xi$ are defined as the modes which are eigenfunctions of the Lie derivative in the $\xi$ direction. Boundary conditions are imposed by defining the vector with respect to which the modes have a positive frequency. In a dynamic collapse, the ingoing modes generally originate from an emitter in the Minkowski spacetime, at past null infinity. These are replicated by defining the ingoing modes to be positive with respect to the time like Killing vector $\partial/\partial t$ at past null infinity. So also, the outgoing modes are defined to have a positive frequency with respect to the Killing vector $\partial/\partial U$ on the past horizon.
@@ II.3 Construction of the Unruh State
This formulation, however, is not entirely valid to model a dynamic collapse the RNdS metric, which is de Sitter, and not Minkowskian at past null infinty. Moreover, the cosmological horizon acts as a causal separation, so that a light ray originating from beyond will not be able to cross the horizon. The ingoing modes seen by an observer inside the cosmological horizon, must therefore be generated by emitters in free fall at the cosmological horizon (and not at infinity, as with other metrics). The boundary conditions must be defined at the past horizon and the cosmological horizon of the black hole.
@@
We will use a coordinate system defined in the RNdS spacetime as follows: The tortoise coordinate is defined as
$dr^{*}=\frac{dr}{\Delta},$ (12)
in terms of which the Eddington-Finklestein null coordinates are
@@
The Unruh state is defined as the positive frequency mode solutions of the massless scalar field equation $\phi_{\omega lm}=0$, which for the radial ingoing and outgoing modes are
$\phi^{in}_{\omega}=\frac{1}{\sqrt{4\pi\omega}}e^{-\iota\omega V_{c}},$ (15a)
$\phi^{out}_{\omega}=\frac{1}{\sqrt{4\pi\omega}}e^{-\iota\omega U_{+}}.$ (15b)
Since modes generated from past null infinity undergo an exponential redshift as they traverse through the collapsing shell of matter, it is the high frequency modes that are of direct relevance to the calculation. Moreover, the high frequency modes are less subject to scattering. The choice of mode solutions as being proportional to the exponent of the proper time implicitly builds in the geometric optics approximation (also known as the high frequency approximation), thereby reducing a scattering problem to a ray tracing problem. We will justify this construct by showing that the proper time of an emitter in free fall at the cosmological horizon is proportional to $V_{c}$, and likewise for an emitter at the event horizon. For convenience, and in accordance with the natural form of the metric, both the emitter and the observer will be designed to start from rest at a point in between the event and cosmological horizons. From that point, which is akin to an unstable equilibrium, they could either fall inwards towards the event horizon or outwards towards the cosmological horizon. It is important to note that a “free falling” emitter at the cosmological horizon, generating ingoing modes, is travelling to a larger spatial radius. The derivative of the outgoing null coordinate is:
@@ III.1 The effective temperature functions
The roots of the horizon function, given in (6) determine the positions of the horizons for a black hole of the determined parameters. The forms of these roots are neither analytically straightforward, nor particularly illuminating. A comprehensive description of the roots may be found in [22]. To reduce the number of independent parameters in the horizon function, as in (6), we define the de Sitter radius $L_{dS}$, related as $L_{dS}^{2}=\frac{3}{\Lambda}$. The mass and charge parameters ($m$ and $q$), are scaled by the de Sitter radius, and the new independent variables $M=\frac{m}{L_{dS}}$ and $Q=\frac{q}{L_{dS}}$ are employed in all the subsequent calculations.
Hereon, a more general form of the horizon function will be employed:
@@
where $r_{-},\ r_{+}$ and $r_{c}$ are the inner (Cauchy), event and cosmological horizons respectively. The fourth root, $\tilde{r}$ is negative and is not of great physical significance. The surface gravities take the form
@@
In addition, the apparent surface gravity at the position of the observer, given by $r_{ob}$, can be defined as
$\varkappa_{ob}=\frac{1}{2}\frac{d\Delta}{dr}\bigg{|}_{r_{ob}}$ (22)
For radial trajectories, the effective potential takes the form $\frac{1}{2}(\Delta-1)$ [23]. The horizon function therefore has a geometric construct similar to that of the potential function associated with the spacetime. Geometrically, the horizon function is always positive in between the event horizon and the cosmological horizon, which requires the existence of a maximum in the horizon function in the region—corresponding to an unstable equilibrium. The observer and emitter can be defined to be at rest at this point of maximum potential, so that their specific energy can be taken as the square root of the value of the horizon function at this point, and will remain constant as they traverse the geodesic. We will here on refer to this point as the rest point ($r_{0}$). An observer starting out from the rest point can either fall in towards the black hole, or outwards towards the cosmological horizon and beyond. There will therefore be two categories of emitters/observers—the infallers and the outfallers. Ingoing modes are generated by a family of emitters (freely outfalling) at the cosmological horizon, outgoing modes are similarly generated by a family of emitters in free fall (infalling) at the event horizon.
@@
$\begin{split}\kappa&=-\frac{d}{d\tau_{ob}}\ln\left(\frac{\omega_{ob}}{\omega_{em}}\right)\\ &=\mp\frac{1}{2}\frac{\omega_{ob}}{\omega_{0}}\left(\frac{d\Delta_{ob}}{dr_{ob}}-\frac{d\Delta_{em}}{dr_{em}}\right),\end{split}$ (23)
@@
$\kappa_{i}^{+}=-\frac{E+\sqrt{E^{2}-\Delta}}{\Delta L_{dS}}\left(\varkappa_{ob}-\varkappa_{+}\right),$ (24a)
$\kappa_{i}^{c}=\frac{E-\sqrt{E^{2}-\Delta}}{\Delta L_{dS}}\left(\varkappa_{ob}-\varkappa_{c}\right),$ (24b)
$\kappa_{o}^{+}=-\frac{E-\sqrt{E^{2}-\Delta}}{\Delta L_{dS}}\left(\varkappa_{ob}-\varkappa_{+}\right),$ (24c)
$\kappa_{o}^{c}=\frac{E+\sqrt{E^{2}-\Delta}}{\Delta L_{dS}}\left(\varkappa_{ob}-\varkappa_{c}\right).$ (24d)
The superscript +(c) refers to the modes originating from the event(cosmological) horizons i.e. the outgoing(ingoing) modes. The subscript i(o) refers to an ingoing(outgoing) observer.
A preliminary analytical examination of the functions reveals that both $\kappa_{o}^{+}$ and $\kappa_{o}^{c}$ tend to constants at the cosmological horizon:
@@
For a black hole in complete thermal equilibrium (at $\varkappa_{+}=\varkappa_{c}$), the ingoing modes see an infinite blueshift at the inner horizon,by virtue of the specific energy going to zero. For outgoing modes, in the limit as $r\rightarrow r_{-}$, the temperature generally diverges, but has a first order asymptotic form

### v2

# Hawking radiation inside a charged, cosmological black hole

@@ Abstract
We study the effective temperature, as a rate of gravitational redshift, of the Hawking modes perceived by a radially free falling observer at an arbitrary location in a Reissner-Nordström-(anti-)de Sitter [RN(A)dS] spacetime. In particular, the behavior of the modes at the inner horizon, and therein the validity of the strong cosmic censorship conjecture under the effective temperature formalism, has been analyzed across the physically permissible parameter space of the RNdS metric. The modes perceived by observers of both positive and negative specific energies have been taken into consideration. Finally, the behavior of the adiabatic control function has been examined over the position space of the observer to determine the regimes where the effective temperature function yields a Planckian spectrum of thermal radiation.
@@ I Introduction
The formulation relevant to this paper is based upon the result that an exponential relation between the affine parameters (u,U) of past and future null infinities results in Hawking radiation [4, 5, 6]. A function $\kappa(\textit{u})$ is defined, dependent on the parametrized relationship U(u), that primarily governs the peeling properties of the null geodesics [7]. It has been shown that given an exponential relation between u and U, and given that the adiabatic condition [Eq. (19)] holds, a thermal flux will be generated, whose temperature is proportional to $\kappa(\textit{u})$. In the limit of a static black hole, $\kappa$ becomes equal to the surface gravity $\varkappa_{+}$ of the event horizon. It is possible to recast the definition of the effective temperature $\kappa$ as the rate of exponential redshift of the Hawking modes. In the case of a collapsing mass, the modes experience a blueshift as they approach the mass and are subject to a gravitational redshift as they climb out. Another explanation that well complements this rather nonintuitive formalism notes that the vacuum state in which the thermal radiance is observed is related to the local Minkowski frame of the observer via an exponential scale transformation [8]. This inherently draws upon the analogous relation between the Unruh effect and the Hawking radiation—an accelerated observer in the Minkowski spacetime, and observer at a constant distance from the black hole will both detect a flux of particles [9].
@@
Of significant consequence in the study of an RNdS black hole, is the existence of a cosmological horizon (at radius $r_{c}$), a boundary that causally separates an observer inside ($r<r_{c}$) from the external universe ($r>r_{c}$). It is characteristically very similar to the event horizon, following the same classical and thermodynamic laws, therein radiating particles with a thermal spectrum [1]. A rigorous study on the thermodynamic properties of the cosmological horizon was presented by Hawking and Gibbons [18], extending to particle creation in de Sitter spaces. While the matter of particle creation in de Sitter space had for long been a topic of study, it had always been viewed in an observer-independent frame. If particle creation were observer independent, and the invariance of the de Sitter group [19, 20, 21] were to hold, an observer Lorentz-boosted into a new frame will see exactly the same spectrum as they would have before the boost. This cannot be true unless the rate of particle production was either zero or infinity, zero being the favored choice of the two. But no particle production would imply no thermal radiation emanating from the cosmological horizon. However it was shown that an observer accelerating in Minkowski spacetime will detect particle production, indicating that particle production must be dependent on the worldline of the observer. An observer moving on a timelike geodesic will detect an isotropic thermal radiation of temperature $T_{H}=\frac{\varkappa_{c}\hbar}{2\pi ck_{B}}$, where $\varkappa_{c}$ is the surface gravity of the cosmological horizon. Once again, the invariance of the de Sitter group requires that any other observer also on a timelike geodesic detect the same radiation—indicating that particle production is in fact observer dependent. We will arrive at the same conclusion, albeit via a different approach. Both the outgoing and the ingoing modes tend to the same isotropic temperature asymptotically far from the black hole, independent of the observer’s specific energy. We will further extend the analysis to negative observer energies and a negative cosmological constant.
@@ II Formalism
II.1 Geodesics in the RNdS geometry
@@
The temporal component of the four velocity of a radially free-falling observer is
@@
Here, we have defined the specific energy of the observer as being the covariant temporal component of the observer’s four velocity. This quantity is constant, and in a way, characterizes the geodesic.
The components for the four velocity of an observer ($u_{\mu}$) and of the null particle ($k^{\mu}$) are found using the standard conditions
@@
The frequency of the null particle as seen by the observer may be obtained by transforming the wave four vector of the null particle into the frame of the observer
@@
where the upper (lower) signs correspond to outgoing (ingoing) modes respectively. $E$ is the specific energy of the observer, which remains constant along the natural geodesic. It was defined earlier as the negative temporal component of the velocity four vector. The value of the specific energy can be easily calculated as the square root of the horizon function at a point where $u_{r}=0$. We can conveniently choose (as we subsequently will) the rest point of the massive particle to be at the point where the horizon function is at a maximum.
@@
We will define the specific energy to be positive for an observer who starts out from rest at a point between the event and cosmological horizons. An infalling (outfalling) observer may be defined as one who travels to a smaller (larger) spatial radius by virtue of the natural geodesic trajectory. An ingoer (outgoer) is defined by their intended direction of motion. The distinction may not appear to be of great significance in a nontrapping region, where an observer can accelerate enough to be able to oppose the natural geodesic. In a trapping region, such as inside the event horizon, it is not possible to be able to travel to a larger spatial radius, and an attempted outward acceleration will induce the observer into a state of negative energy. An observer beyond the event horizon ($r<r_{+}$), who is infalling and outgoing, will have a specific energy $E<0$. So also, an observer beyond the cosmological horizon who is outfalling but ingoing will also have a negative specific energy. Note that the modes are defined as being either outgoing or ingoing, a state determined by the initial conditions at past null infinity. Outgoing modes inside the event horizon will remain outward directed, but will travel to a smaller spatial radius, eventually undergoing a divergence at the inner horizon.
II.2 Construction of the Unruh quantum state
In order to define the vacuum state of the quantum field in the spacetime geometry defined by Eq. (1), the massless scalar wave equation, $\phi_{\omega lm}=0$, may be solved. The field may be decomposed into a set of orthonormal positive and negative frequency modes
[…]
The modes can be further separated into radial and angular components, of which the radial component, $f_{\omega l}$ must satisfy
$\frac{\partial^{2}f_{\omega l}}{\partial r^{*2}}-\frac{\partial^{2}f_{\omega l}}{\partial t^{2}}=\Delta\left(\frac{l(l+1)}{r^{2}}+\frac{1}{r}\frac{d\Delta}{dr}\right)f_{\omega l}.$ (9)
The tortoise coordinate $r^{*}$ is defined as
$dr^{*}=\frac{dr}{\Delta}.$ (10)
The solutions to $f_{\omega l}$ yields two sets of waves—one coming in from infinity, the other emanating from the past horizon of the black hole, that is, the ingoing and outgoing modes respectively.
The quantum field’s vacuum state is formally defined by canonical quantization of the ladder operators $a_{\omega lm}$ of Eq. (8), which practically manifests itself as the choice of past boundary conditions for the mode solutions $f_{\omega l}$ in Eq. (9). The asymptotic past of a realistic black hole is flat everywhere, so that the vacuum state can be unambiguously defined by initializing a set of ingoing modes $f_{\omega l}=e^{-i\omega(r^{*}+t)}$ at infinity and propagating these modes through the collapsing matter to obtain outgoing modes. However, the RNdS metric used throughout this paper is eternal and static, and therefore it contains a horizon at $r_{+}$ in its asymptotic past boundary. The Unruh quantum state [9] solves this problem by initializing a set of outgoing modes at the past horizon so that the resulting quantum modes throughout the spacetime are equivalent to those obtained by a collapse model taken far enough into the past.
A minor subtlety with the choice of the Unruh state to characterize the spacetime’s semiclassical behavior is that the Unruh modes diverge everywhere along the Cauchy horizon, whereas in some shell collapse models [22], the state can be made to remain regular as it passes through the outgoing portion of the inner horizon. However, regular evolution through the inner horizon is not expected to be generic at either the classical or semiclassical level as a result of the perturbation-induced mass inflation instability [23]. Thus, the Unruh state used here should be a reasonable approximation for the semiclassical late-time behavior of realistic black holes.
“Emitters” will hereon represent the frame with respect to which the field is quantized. In an asymptotically Minkowskian spacetime, the ingoing modes are generated from a family of emitters in free fall at past null infinity. The modes are therefore positive frequency with respect to the proper time of a freely falling emitter in Minkowski spacetime. Similarly, the outgoing modes are generated from a family of freely falling emitters at the past horizon, and therefore defined to have a positive frequency with respect to the proper time of this family of emitters. “Observers”, a term used previously, are simply representatives of an entity that has the ability to detect the radiation at arbitrary locations.
@@ II.2 Construction of the Unruh quantum state
The RNdS metric, however, is de Sitter, and not Minkowskian at past null infinity. Moreover, the cosmological horizon acts as a causal separation, so that a light ray originating from beyond will not be able to cross the horizon. The ingoing modes seen by an observer inside the cosmological horizon, must therefore be generated by emitters in free fall at the cosmological horizon (and not at infinity, as with other metrics). The boundary conditions must be defined at the past horizon and the cosmological horizon of the black hole.
@@
We will use a coordinate system defined in the RNdS spacetime as follows: In terms of the tortoise coordinate, as defined in Eq. (10), the Eddington-Finklestein null coordinates are
@@
In finding the normalized mode solutions for a spherically symmetric, static black hole, the angular components can be safely suppressed [4]. This permits us to express the modes in a 1+1D spacetime. Additionally, the Polyakov approximation that permits a dimensional reduction of the quantum stress energy tensor for spherically symmetric, static and conformal metric, to a two-dimensional spacetime does not withhold any significant properties of the modes, in the geometric optics limit (which has already been assumed) [23]. In this light, we may express the mode solutions in a 1+1D spacetime as
$f^{in}_{\omega}=\frac{1}{\sqrt{4\pi\omega}}e^{-\iota\omega V_{c}},$ (13a)
$f^{out}_{\omega}=\frac{1}{\sqrt{4\pi\omega}}e^{-\iota\omega U_{+}}.$ (13b)
Since modes generated from past null infinity undergo an exponential redshift as they traverse through the collapsing shell of matter, it is the high frequency modes that are of direct relevance to the calculation. Moreover, the high frequency modes are less subject to scattering. The choice of mode solutions as being proportional to the exponent of the proper time implicitly builds in the geometric optics approximation (also known as the high frequency approximation), thereby reducing a scattering problem to a ray tracing problem. We will justify this construct by showing that the proper time of an emitter in free fall at the cosmological horizon is proportional to $V_{c}$, and likewise for an emitter at the event horizon. For convenience, and in accordance with the natural form of the metric, both the emitter and the observer will be designed to start from rest at a point in between the event and cosmological horizons. From that point, which is akin to an unstable equilibrium, they could either fall inward toward the event horizon or outward toward the cosmological horizon. It is important to note that a “free falling” emitter at the cosmological horizon, generating ingoing modes, is traveling to a larger spatial radius. The derivative of the outgoing null coordinate is:
@@
II.3 The effective temperature formalism
Having defined both the geometry and the quantum state coupled to it, we may now offer a precise formulation of the effective temperature function.
The objective, as previously mentioned, is to generalize the calculation of the original Hawking temperature to arbitrary observers. The perception of thermal radiation, as shown in [8], arises from an exponential relation between the affine parameters of the emitter and the observer. Such an exponential relation is known to exist for a freely falling observer-emitter pair in a black hole metric (in which the observer detects a Hawking like radiation), or in a Rindler-Minkowski frame (in which the observer in the Rindler space sees an Unruh radiation).
If a Hawking flux is to be perceived, one can therefore define an exponential relation between the two affine parameters. Consider an emitter following a trajectory characterized by affine parameter $U$, and an observer following a timelike geodesic characterized by affine parameter $u$. The worldlines of the emitter and the observer are connected by a null ray following a geodesic path given by $U(u)$. Define a parametric relation $U(u)$ [4], via a function $\kappa(u)$:
An integration of the above equation, to obtain $U(u)$ shows that this parametrization does indeed characterize an exponential relation. Physically, the adiabatic condition requires that around any given curve characterized by $u_{*}$, the value of the function $\kappa_{*}\equiv\kappa(u_{*})$ remains approximately constant. Mathematically, the constraint can be given as
$\left|\frac{1}{\kappa_{*}^{2}}\frac{d\kappa_{*}}{du}\right|\ll 1.$ (19)
$T_{H}=\frac{\kappa(u_{*})}{2\pi}.$ (20)
Having scaled the constants, the effective temperature at any point can be well given by $\kappa(u_{*})$. Of particular note is that this formalism is independent of the position $u_{*}$ of the observer. $\kappa(u)$ by definition, is the parametrized relation between the affine parameters of the emitter and the observer, describing the peeling properties of the null geodesics. It is on this quantity that the Hawking temperature is dependent, rather than the surface gravity itself. In the limit of a static black hole, i.e., as the observer approaches future infinity, $\kappa(u)$ equals the surface gravity, reducing the expression to the well established Hawking temperature. This formulation Eq. (20) thereby permits a calculation of the effective temperature perceived by an observer at any point in spacetime.
The affine parameter that characterizes the trajectory of any particle is the proper time. Therefore, the affine parameters u and U may be relabeled with the proper times $\tau_{ob}$ and ${\tau_{em}}$ of the observer and emitter respectively. The effective temperature function Eq. (18) can be written as
In this form, the effective temperature function is seen to be the rate of gravitational redshift. It is this form of the effective temperature function that will be relevant to the subsequent calculations.
@@ III.1 The effective temperature functions
The roots of the horizon function, given in Eq. (2) determine the positions of the horizons for a black hole of the determined parameters. The forms of these roots are neither analytically straightforward, nor particularly illuminating. A comprehensive description of the roots may be found in [24]. To reduce the number of independent parameters in the horizon function, as in Eq. (2), we define the de Sitter radius $L_{dS}$, related as $L_{dS}^{2}=\frac{3}{\Lambda}$. The mass and charge parameters ($m$ and $q$), are scaled by the de Sitter radius, and the new independent variables $M=\frac{m}{L_{dS}}$ and $Q=\frac{q}{L_{dS}}$ are employed in all the subsequent calculations.
Hereon, a more general form of the horizon function will be employed
@@
where $r_{-},\ r_{+}$ and $r_{c}$ are the inner (Cauchy), event, and cosmological horizons, respectively. The fourth root, $\tilde{r}$ is negative and is not of great physical significance. The surface gravities take the form
@@
Surface gravities arise from the geodesic equations on a Killing horizon, and it is on this basis that they are defined. One may also define a “generalized surface gravity” [25] anywhere in the spacetime for an observer on an imaginary surface with radius $r_{ob}$ as
$\varkappa(r_{ob})=\frac{1}{2}\frac{d\Delta}{dr}\bigg{|}_{r_{ob}}.$ (24)
For radial trajectories, the effective potential takes the form $\frac{1}{2}(\Delta-1)$ [26]. The horizon function therefore has a geometric construct similar to that of the potential function associated with the spacetime. Geometrically, the horizon function is always positive in between the event horizon and the cosmological horizon, which requires the existence of a maximum in the horizon function in the region—corresponding to an unstable equilibrium. The observer and emitter can be defined to be at rest at this point of maximum potential, so that their specific energy can be taken as the square root of the value of the horizon function at this point, and will remain constant as they traverse the geodesic. We will here on refer to this point as the rest point ($r_{0}$). An observer starting out from the rest point can either fall in toward the black hole, or outward toward the cosmological horizon and beyond. There will therefore be two categories of emitters/observers—the infallers and the outfallers. Ingoing modes are generated by a family of emitters (freely outfalling) at the past cosmological horizon, outgoing modes are similarly generated by a family of emitters in free fall (infalling) at the past event horizon.
@@
$\begin{split}\kappa&=-\frac{d}{d\tau_{ob}}\ln\left(\frac{\omega_{ob}}{\omega_{em}}\right)\\ &=\mp\frac{1}{2}\frac{\omega_{ob}}{\omega_{0}}\left(\frac{d\Delta}{dr}\bigg{|}_{r_{ob}}-\frac{d\Delta}{dr}\bigg{|}_{r_{em}}\right),\end{split}$ (25)
@@
$\kappa_{i}^{+}(r_{ob})=-\frac{E+\sqrt{E^{2}-\Delta(r_{ob})}}{\Delta(r_{ob})L_{dS}}\left(\varkappa(r_{ob})-\varkappa_{+}\right),$ (26a)
$\kappa_{i}^{c}(r_{ob})=\frac{E-\sqrt{E^{2}-\Delta(r_{ob})}}{\Delta(r_{ob})L_{dS}}\left(\varkappa(r_{ob})-\varkappa_{c}\right),$ (26b)
$\kappa_{o}^{+}(r_{ob})=-\frac{E-\sqrt{E^{2}-\Delta(r_{ob})}}{\Delta(r_{ob})L_{dS}}\left(\varkappa(r_{ob})-\varkappa_{+}\right),$ (26c)
$\kappa_{o}^{c}(r_{ob})=\frac{E+\sqrt{E^{2}-\Delta(r_{ob})}}{\Delta(r_{ob})L_{dS}}\left(\varkappa(r_{ob})-\varkappa_{c}\right).$ (26d)
The superscript + (c) refers to the modes originating from the past event (cosmological) horizon, i.e. the outgoing (ingoing) modes. The subscript i (o) refers to an ingoing (outgoing) observer. Fig. 1 shows the trajectories of the observer and the emitter, as well as the modes corresponding to the four effective temperatures in Eqs. (26). This being the Unruh state, the emitter is asymptotically close to the boundaries, and the observer can be either ingoing or outgoing from the point of maximum potential.
Figure 1: The orange (purple) lines represent the trajectories of the observer (emitter), and the solid (dashed) lines correspond to an ingoing (outgoing) trajectory. The red (yellow) and the blue (green) arrows represent the outgoing and ingoing modes respectively seen by an ingoing (outgoing) observer.
Though these effective temperatures may at first glance appear to diverge at the cosmological horizon due to the vanishing horizon function $\Delta(r_{ob})$ in the denominator of Eqs. (26), they nonetheless remain finite, since the numerator term $E-\sqrt{E^{2}-\Delta(r_{ob})}$ also vanishes as $\mathcal{O}(\Delta)$ for outgoing modes, while for ingoing modes, the term $\varkappa(r_{ob})-\varkappa_{c}$ cancels out the blueshift divergence. A preliminary analytical examination of the functions reveals that both $\kappa_{o}^{+}$ and $\kappa_{o}^{c}$ tend to constants at the cosmological horizon
@@
However, for outgoing modes, in the limit as $r\rightarrow r_{-}$, the effective temperature generally diverges, since the state-dependent surface gravity term $\varkappa(r_{ob})-\varkappa_{+}$ in Eq. (26a) generally does not vanish to cancel out the diverging blueshift term $(E-\sqrt{E^{2}-\Delta})/\Delta$. The effective temperature has first-order asymptotic form